Dear Sir,
Please help me in giving the solution to the foll.problem as I am a chemical engineer in doing simulation softwares
Please reply how to determine the value of THETTA in the foll.problem
((THETTA) - SIN(THETTA)) = 0.5268
Regards,
Prem.

Answer by khwang(438)

Since there is no way to get the exact solution the equation x - sin x = c (constant).
We only can only use series approximations by Taylor series:
sin x ~ x -x^3/6,
To solve x - sin x = 0.5268
That is, x^3/6 = 0.5268,
so, x = (0.5268 * 6)^1/3 = 0.1536 (approximated value)
Kenny